# explExample

### v2

@@ 1 Introduction
Various attempts have been made to obtain the reduced complexity for convection-dominated problems by incorporating the structure of the transport. To provide a possibly incomplete list of proposed approaches that are the most relevant to this work: template-fitting [74], method of freezing [4, 59], shock reconstruction [16], approximated Lax-Pairs [28], advection modes [40], transported snapshot interpolation (TSI) [83], shifted proper orthogonal decomposition (sPOD) [67], calibrated manifolds [9], Lagrangian basis method [54], transport reversal [70], registration methods [78], Wasserstein barycenters [24] or quadratic manifolds [26]. Some of these proposed methods can be viewed as dynamical low-rank approximations [43, 75].
@@
The compositional reduced models, while achieving the two properties required of classical reduced models, have not yet demonstrated an ability to handle shock propagation in their full generality; they were tested in restricted settings without shock interaction. This is a serious limitation, since for many PDEs arising from fluid dynamics, shock formation and collision are defining features. For example, a typical entropy solution to the Burgers’ equation is shown in Figure 1. Shocks can form at various times, travel at varying speeds, and merge; new techniques are required to handle these challenges. For example, a computational approach called Front Transport Reduction (FTR) employs level sets to represent the changes in the shock topology [44].
@@
The third property has important practical implications: Despite the involved spatial approximation, the temporal dependence of the weight parameters is linear. This suggests that this model has the potential to approximate complicated solutions while maintaining regular dependence on time. Such a feature would enable one to computationally integrate the representation in time, regardless of the lack of regularity in the spatial variable due to shocks. We briefly mention here a close relation between regular-in-time approximation introduced here and classical numerical methods exploiting similar properties, e.g. Large Time Step methods [49] or Averaged Multivalued Solutions [8].
@@
Our construction here is a similarly parametrized family of INRs with low-rank architecture we call the Low-Rank Neural Representation (LRNR). We use the term LRNR to distinguish the technical definition of the architecture alone, as separate from the machine learning context, and to contrast with RDNs [72] which took on a different approach to approximating the weak solution. Readers familiar with the LR-PINNs may view LRNR as a version of the architecture used in Low-Rank PINNs [13], keeping in mind that the time variable is treated as a parameter here rather than an input. We remark that LRNRs may be widely useful in general data-driven settings to produce a dimension reduction for data sets that are suitably described by INRs, independent of the PINNs framework.
The trained LRNRs can serve as a reduced model approximation of the family of solutions to the parametrized PDEs: A rapid approximation is achieved due to the small degrees of freedom, much like the rapid computations performed during the online phase for classical reduced models. Moreover, it is not difficult to see that a reduction in online complexity can be achieved by utilizing the low rank structure in the weights and biases,  [^1] During the review of this work, a concrete result demonstrating this observation has appeared in [12]. much like the classical reduced models [38]. This is a property LRNR inherits by being a direct generalization of the compositional reduced model utilizing transported subspaces [71].
@@
  • Section 5. We develop a LRNR approximation to the entropy solutions by building on the new representation of the entropy solution developed in Section 2 and the inverse-bias trick from Section 4. We show that general entropy solutions can be approximated by a LRNR with five layers and rank two. This construction establishes a general low dimensional representation of the entropy solution, enabling an efficient representation of arbitrary shock wave interactions.
@@ 2.1 Entropy solutions
(2.3) $\overline{\mathcal{U}}:=\{v\in\mathcal{U}\cap P_{0}(\Omega_{x})\mid v\text{ has finitely many jumps }\}$
@@
(2.4) $G(X,t):=\left\{\begin{aligned} F^{\prime}(u(X,t))&\quad\text{ if }\llbracket u(X,t)\rrbracket=0,\\ \frac{\llbracket F(u(X,t))\rrbracket}{\llbracket u(X,t)\rrbracket}&\quad\text{if }\llbracket u(X,t)\rrbracket\neq 0.\\ \end{aligned}\right.$
The function $G$ yields the velocity of the classical characteristic curves if the point $(X,t)\in\Omega$ lies in a smooth part of $u$, or yields the shock speed via the Rankine-Hugoniot jump conditions if it lies on a shock curve where $u$ is discontinuous.
@@
(2.6) $G(X,x,t):=\left\{\begin{aligned} F^{\prime}(u_{0}(x))&\quad\text{ if }\llbracket u(X,t)\rrbracket=0,\\ \frac{\llbracket F(u(X,t))\rrbracket}{\llbracket u(X,t)\rrbracket}&\quad\text{if }\llbracket u(X,t)\rrbracket\neq 0.\\ \end{aligned}\right.$
@@
Since our problem has no source terms (2.1), it is constant along the characteristic curves. So we may rewrite (2.5) as follows,
@@
(2.8) :=$\mathcal{R}:=\left\{(x,t)\in\Omega\mid x\in\Omega_{x}\smallsetminus X(\Omega_{x},t)\right\}.$
@@
(2.11) $u(x,t)=\left\{\begin{aligned} u_{0}(X^{+}(x,t))\quad&\text{ if }(x,t)\in\Omega\smallsetminus\mathcal{R},\\ (F^{\prime})^{-1}\left(\xi(x,t)\right)\quad&\text{ if }(x,t)\in\mathcal{R}.\end{aligned}\right.$
@@ 2.2 Rarefied characteristics
Observe that the set $\overline{\xi}(\mathcal{R})$ is finite, since there are only finitely many jumps in the piecewise constant initial condition $u_{0}\in\overline{\mathcal{U}}$ from which the rarefaction fans originate. Let $n_{\mathcal{R}}:=|\overline{\xi}(\mathcal{R})|$ then we order and enumerate these by
@@
(2.15) $\nu_{k}^{+}:=\sup_{(x,t)\in\mathcal{R}_{k}}\frac{x-\xi_{k}}{t},\qquad\nu_{k}^{-}:=\inf_{(x,t)\in\mathcal{R}_{k}}\frac{x-\xi_{k}}{t}.$
@@
(2.19) $\left\lvert\Gamma_{x}\right\rvert=\sum_{k=1}^{n_{\mathcal{R}}}|\Gamma_{k}|=\sum_{k=1}^{n_{\mathcal{R}}}\varepsilon_{k}=\left\lvert u_{0}\right\rvert_{PV(\Omega_{x})}.$
Next we extend the initial condition $u_{0}$ to $\widehat{\Omega}_{x}$ by letting $\hat{u}_{0}:\widehat{\Omega}_{x}\to\mathbb{R}$ be given as
@@
Intuitively speaking, one obtains $\hat{u}_{0}$ from $u_{0}$ by replacing positive jumps with continuous slopes, prescribed the similarity solution. See Figure 2 for an illustration comparing $u_{0}$ and $\hat{u}_{0}$.
Figure 2: An example of an extended initial condition $\hat{u}_{0}$ of $u$. The newly assigned values of $\hat{u}_{0}$ in the interval $\Gamma_{1}$ are given by (2.20).
@@
(2.24) $\widehat{G}(\widehat{X},x,t):=\left\{\begin{aligned} F^{\prime}(\hat{u}_{0}(x))&\quad\text{ if }\llbracket u(\widehat{X},t)\rrbracket=0,\\ \frac{\llbracket F(u(\widehat{X},t))\rrbracket}{\llbracket u(\widehat{X},t)\rrbracket}&\quad\text{ if }\llbracket u(\widehat{X},t)\rrbracket\neq 0.\\ \end{aligned}\right.$
@@
See Figure 3 for an illustration showing $\hat{I}$ and its role in relating $u_{0}$ with $\hat{u}_{0}$ (2.22). A diagram depicting the evolution of $\widehat{X}$ is shown in Figure 4.
@@
Let us define the rarefied characteristics with multivalued inverse,
@@ 2.3 Shock time function
Let us define the parametrized set shock set $\Upsilon_{t}$ for $t\in\Omega_{t}$ to contain all $x\in\widehat{\Omega}_{x}$ such that the preimage of the singleton $\{x\}$ given by $\widehat{X}^{-1}(\widehat{X}(\{x\},\,t),\,t)$ is non-unique. That is,
@@
Note that $\Upsilon_{t}$ is closed for all $t\in\Omega_{t}$; $\widehat{X}(\cdot,t)$ is non-increasing and continuous, implying that $\Upsilon_{t}$ is a union of disjoint closed intervals. It follows that $\lambda$ is continuous: a closed set $D$ lying in $\Omega_{t}$ is compact, and moreover $\Upsilon_{t}$ is increasing in $t$, hence we have $\lambda^{-1}(D)=\Upsilon_{\max D}$, which is closed.
@@
It is straightforward to see that $\widehat{X}$ is related to $\widehat{X}_{0}$ by the relation,
@@ 2.4 Compositional form
In the previous section, we have simplified the expression of the entropy solution from (2.11) to (B.1), by incorporating the evolution of the rarefaction waves into the characteristic curves in (2.22). Next, we simplify how the shock is represented in $\widehat{X}$.
In the classical regime, the characteristic curves (2.5) are represented by a linear superposition,
@@
We wish to rewrite $\widehat{X}(x,t)$ in a form with similar linear superpositions, modulo compositions and left-inverses. Note that one can show the family of functions $\{\widehat{X}(\cdot,t)\mid t\in\Omega_{t}\}$ no longer has a fast decaying Kolmogorov width, except in special cases with simple initial data $u_{0}$. Consider the stationary shock example shown in Figure 5. The shock is represented by the constant region $\Upsilon_{t}$ in $\widehat{X}$, which is expanding over time. The jump in the derivative in $\widehat{X}$ travels to the left and the right, which cannot be efficiently represented by the superposition of a few basis functions.
Figure 5: A stationary shock example. Due to the jump discontinuity in the initial condition $\hat{u}_{0}$ located at $x=0.5$, the shock forms immediately at initial time. This shock is represented as the constant part of the function $\widehat{X}$ in $\Upsilon_{t}$, and the left-inverse of the rarefied characteristic $\widehat{X}^{+}$ has a jump discontinuity there. As time evolves, the characteristics merge into the shock, so $\Upsilon_{t}$ where $\widehat{X}$ is constant correspondingly expands over time. Consequently, the kinks in $\widehat{X}$ at the left and right endpoints of $\Upsilon_{t}$ travels to the left and to the right, respectively. (For a plot of the evolution of the rarefied characteristics, see Figure 7 below.)
Our key idea is to cut out the portion of the domain where $\widehat{X}_{0}(\cdot,t)$ represents the shock, by adding the characteristic function of $\Upsilon_{t}$ to it. Let us denote the threshold function by
(2.31) $\rho:=\mathbf{1}_{\mathbb{R}_{+}}.$
We call the modified curves the relief characteristic curves, given by
@@
Once the left-inverse is applied to $\widecheck{X}$, the resulting map $\widecheck{X}^{+}:\Omega_{x}\to\widehat{\Omega}_{x}$ is constant at points in $\Upsilon_{t}$, because the term $\rho(t-\lambda(x))$ introduces the constant region and removes the portion of $\widehat{X}$ which lies in $\Upsilon_{t}$ by adding a sufficiently large constant to $\widehat{X}_{0}$ whenever it is evaluted at a point in $\Upsilon_{t}$. See Figure 6 for an illustration.
@@
In this representation, the time-dependence is in the coefficients of linear combinations within each composition, and the dependence of the coefficients with respect to $t$ is linear, in a similar manner as the classical case. To elaborate, the first linear combination in (2.34) is $t\varphi_{1}-\varphi_{2}$ with two coefficient values $t$ and $-1$. Only the first coefficient depends on $t$ and the dependence is linear.
In the compositional form, the left-inverse $(\cdot)^{+}$ persists, leaving it an implicit form rather than an explicit one. When we consider the discretization of this form of the entropy solution, we will make use of a trick we call the inverse-bias trick that will allow us to avoid any direct evaluation of the left-inverse (Section 4.1).
@@
Figure 6: A diagram depicting the relief characteristics $\widecheck{X}$ after shock formation time $t_{S}$ is shown, along with the rarefied characteristics with multivalued inverse $\widehat{X}_{0}$. The relief characteristics use a different representation of the shock compared to the relief characteristics $\widehat{X}$, one that uses the addition of a box function $1_{\Upsilon_{t}}$. One observes that left-inverse $\widecheck{X}^{+}$ has a jump at $s(t)$, and this jump is identical to the corresponding one for $\widehat{X}^{+}$.
@@ 3 Low Rank Neural Representation
(3.2) [$\ \phi_{1}(z)&\phi_{2}(z)&\phi_{3}(z)\\ \mid&\mid&\mid\end{bmatrix}\begin{bmatrix}1\\ t\\ 1\end{bmatrix}\quad\in\operatorname{span}\{\phi_{1},\phi_{2},\phi_{3}\}$
where $z=(x,t\varphi_{1}(x)-\varphi_{2}(x))^{T}$ is the output from the previous step. This expression is superposition of a few vectors, if one views the function values $\phi_{i}(z)$ as column vectors of infinite length. In addition, note that the individual the coefficients $[1,t,1]^{T}$ has either a linear dependence on $t$ or is constant.
This structure of repeated compositions closely resembles the repeated compositions in a neural network. The component that requires a close scrutiny is the left-inverse $(\cdot)^{+}$ and its treatment will be a key discussion in the coming sections (Sections 4.1 and 5.1). For the moment, let us focus on establishing an analogy with feedforward neural networks. We begin by introducing notations and definitions of a neural network. Our notations are similar to those in common use (see, e.g. [22, 3], for similar definitions). Throughout, the superscript $\ell$ will index the layer number.
@@
We will use the shorthand $h(W,B):=h(\ \cdot\ ;W,B)$.
@@
$=\gamma_{1}^{3}\begin{bmatrix}\cdots\cdots\end{bmatrix}\sigma\odot\,\Bigg{(}\underbrace{\gamma_{1}^{2}\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}{\color[rgb]{.5,.5,.5}\begin{bmatrix}\cdots\cdots\end{bmatrix}}}_{W^{2}}{\color[rgb]{.5,.5,.5}\sigma\odot\,\Bigg{(}\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}}\begin{bmatrix}x\end{bmatrix}{\color[rgb]{.5,.5,.5}+\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}\Bigg{)}+}\begin{bmatrix}\vdots\\ \vdots\end{bmatrix}{\color[rgb]{.5,.5,.5}[\cdot]}\Bigg{)}+[\cdot].$
@@
The coefficients $\{\gamma^{\ell}_{i},\theta^{\ell}_{i}\}$ correspond to the coefficients $\{1,t\}$ used in the linear combinations at each compositional step in (3.1). One notes that the latter coefficients depend on the parameter $t$ (the time variable), and the role of this parameter is different from either the input $x$ or the fixed weights and biases $\{U_{i}^{\ell},V_{i}^{\ell}\}$ prescribing the linear spaces the weights and biases belong to. It is therefore natural to view the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ as distinct external parameters; a specific set of coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ correspond to a specific neural network belonging to the family of neural networks whose weights and biases satisfy (3.10). The dependence on $t$ in (3.1) would correspond to the dependence of the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ on $t$. On the other hand, the fixed weights and biases correspond to the functions in the low rank representation that do not depend on $t$, e.g. the functions $\{\varphi_{i},\phi_{j},\psi_{k}\}$ in (3.2).
@@
    (3.14) {$U^{\ell}_{i}&=u^{\ell,L}_{i}\bullet u^{\ell,R}_{i},\\ V^{\ell}_{i}&=v^{\ell,L}_{i}\bullet^{\prime}v^{\ell,R}_{i},\end{aligned}\right.\quad\text{ where }\quad\bullet,\bullet^{\prime}\in\{\otimes,{}_{d}\otimes,\otimes_{d}\},$
@@
    sets of coefficients $C^{\ell},D^{\ell}\subset\mathbb{R}^{r}$ collected in
@@
In summary, a LRNR is a family of feedforward networks that have weights and bias at each individual layer that belong to a fixed linear subspace of dimension at most $r$. The aim of the following sections is to establish the connection between LRNR (Definition 3.3) and the entropy solution (3.1). Throughout this work, our focus is on scalar problems in a single spatial dimension, so the input and output dimensions are one, that is, $M_{0}=M_{L}=1$.

### v4

@@ 1 Introduction
Various attempts were made to obtain the reduced complexity for convection-dominated problems by incorporating the structure of the transport. To provide a possibly incomplete list of proposed approaches that are the most relevant to this work: template-fitting [74], method of freezing [4, 59], shock reconstruction [16], approximated Lax-Pairs [28], advection modes [40], transported snapshot interpolation (TSI) [83], shifted proper orthogonal decomposition (sPOD) [67], calibrated manifolds [9], Lagrangian basis method [54], transport reversal [70], registration methods [78], Wasserstein barycenters [24] or quadratic manifolds [26]. Some of these proposed methods can be viewed as dynamical low-rank approximations [43, 75].
@@
The compositional reduced models, while achieving the two properties required of classical reduced models, have not yet demonstrated an ability to handle shock propagation in their full generality; they were tested in restricted settings without shock interaction. This is a serious limitation, since for many PDEs arising from fluid dynamics, shock formation and collision are defining features. For example, a typical entropy solution to the Burgers’ equation is shown in Fig. 1. Shocks can form at various times, travel at varying speeds, and merge; new techniques are required to handle these challenges. For example, a computational approach called Front Transport Reduction (FTR) employs level sets to represent the changes in the shock topology [44].
@@
The third property has important practical implications: Despite the involved spatial approximation, the temporal dependence of the weight parameters is linear. This suggests that this model has the potential to approximate complicated solutions while maintaining regular dependence on time. Such a feature would enable one to computationally integrate the representation in time, regardless of the lack of regularity in the spatial variable due to shocks. We briefly mention here a close relation between regular-in-time approximation introduced here and classical numerical methods exploiting similar properties, e.g. Large Time Step methods [49] or Averaged Multivalued Solutions [8]: These methods exploit the fact that classical characteristic curves are regular in time, that is, they are straight lines. These methods propagate the characteristics for longer durations than permitted by the Courant-Friedrichs-Lewy condition and then modify them to ensure proper shock propagation. We closely follow this principle of time-regular characteristic evolution with minimal modification. In contrast, standard finite-volume methods alter the characteristic curves at each time step and do not preserve their straightness or temporal regularity [50].
@@
Our construction here is a similarly parametrized family of INRs with low-rank architecture we call the Low-Rank Neural Representation (LRNR). We use the term LRNR to distinguish the technical definition of the architecture alone, as separate from the machine learning context, and to contrast with RDNs [72] which took on a different approach to approximating the weak solution. Readers familiar with the LR-PINNs may view LRNR as a version of the architecture used in Low-Rank PINNs [13], keeping in mind that the time variable is treated as a parameter here rather than an input.
The trained LRNRs can serve as a reduced model approximation of the family of solutions to the parametrized PDEs: A rapid approximation is achieved due to the small number of degrees of freedom, much like the rapid computations performed during the online phase for classical reduced models. Moreover, it is not difficult to see that a reduction in online complexity can be achieved by utilizing the low rank structure in the weights and biases [^1] During the review of this work, a concrete result demonstrating this observation has appeared in [12]. much like the classical reduced models [38]. This is a property LRNR inherits by being a direct generalization of the compositional reduced model utilizing transported subspaces [71].
@@
  • Section 5. We develop a LRNR approximation to the entropy solutions by building on the new representation of the entropy solution developed in Section 2 and the inverse-bias trick from Sec. 4. We show that general entropy solutions can be approximated by a LRNR with five layers and rank two. This construction establishes a general low dimensional representation of the entropy solution, enabling an efficient representation of arbitrary shock wave interactions.
@@ 2.1 Entropy solutions
[…]
@@
[…]
where $\llbracket v(x,t)\rrbracket:=v(x_{+},t)-v(x_{-},t)$ denotes the different between right and left limits of $v(\cdot,t)$ at $x$. The function $G$ yields the velocity of the classical characteristic curves if the point $(X,t)\in\Omega$ lies in a smooth part of $u$, or yields the shock speed via the Rankine-Hugoniot jump conditions if it lies on a shock curve where $u$ is discontinuous.
@@
[…]
@@
Since our problem has no source terms (2.1), the solution is constant along the characteristic curves. So we may rewrite (2.5) as follows,
@@
(2.8) $\mathcal{R}=\left\{(x,t)\in\Omega\mid x\in\Omega_{x}\smallsetminus X(\Omega_{x},t)\right\}.$
@@
[…]
@@ 2.2 Rarefied characteristics
Observe that the set $\overline{\xi}(\mathcal{R})$ is a finite union of single points, since there are only finitely many jumps in the piecewise constant initial condition $u_{0}\in\overline{\mathcal{U}}$ from which the rarefaction fans originate. Let $n_{\mathcal{R}}:=|\overline{\xi}(\mathcal{R})|$ then we order and enumerate these by
@@
(2.15) $\nu_{k}^{+}=\sup_{(x,t)\in\mathcal{R}_{k}}\frac{x-\xi_{k}}{t},\qquad\nu_{k}^{-}=\inf_{(x,t)\in\mathcal{R}_{k}}\frac{x-\xi_{k}}{t}.$
@@
(2.19) $\left\lvert\Gamma_{x}\right\rvert=\sum_{k=1}^{n_{\mathcal{R}}}|\Gamma_{k}|=\sum_{k=1}^{n_{\mathcal{R}}}\varepsilon_{k}=\left\lvert u_{0}\right\rvert_{PV(\Omega_{x})},$
where $\left\lvert\cdot\right\rvert$ denotes the measure of a set. Next we extend the initial condition $u_{0}$ to $\widehat{\Omega}_{x}$ by letting $\hat{u}_{0}:\widehat{\Omega}_{x}\to\mathbb{R}$ be given as
@@
Intuitively speaking, one obtains $\hat{u}_{0}$ from $u_{0}$ by replacing positive jumps with continuous slopes, prescribed the similarity solution. See Fig. 2 for an illustration comparing $u_{0}$ and $\hat{u}_{0}$.
Figure 2: An example of an extended initial condition $\hat{u}_{0}$ of $u_{0}$ in the case of Burgers’ flux. The newly assigned values of $\hat{u}_{0}$ in the interval $\Gamma_{1}$ are given by (2.20). The values are linear due to the choice of the Burgers’ flux.
@@
[…]
@@
See Fig. 3 for an illustration showing $\hat{I}$ and its role in relating $u_{0}$ with $\hat{u}_{0}$ (2.22). A diagram depicting the evolution of $\widehat{X}$ is shown in Fig. 4.
@@
Let us define the rarefied characteristic curves with multivalued inverse,
@@ 2.3 Shock time function
Let us define the parametrized set we call the shock set $\Upsilon_{t}$ for $t\in\Omega_{t}$ to contain all $x\in\widehat{\Omega}_{x}$ such that the preimage of the singleton $\{x\}$ given by $\widehat{X}^{-1}(\widehat{X}(\{x\},\,t),\,t)$ is non-unique. That is,
@@
Note that $\Upsilon_{t}$ is closed for all $t\in\Omega_{t}$; $\widehat{X}(\cdot,t)$ is non-decreasing and continuous, implying that $\Upsilon_{t}$ is a union of disjoint closed intervals. It follows that $\lambda$ is continuous: a closed set $D$ lying in $\Omega_{t}$ is compact, and moreover $\Upsilon_{t}$ is increasing in $t$, hence we have $\lambda^{-1}(D)=\Upsilon_{\max D}$, which is closed.
@@
It is left to the reader to show that $\widehat{X}$ is related to $\widehat{X}_{0}$ by the relation,
@@ 2.4 Compositional form
In the previous section, we have simplified the expression of the entropy solution from (2.11) to (2.29), by incorporating the evolution of the rarefaction waves into the characteristic curves in (2.22). Next, we simplify how the shock is represented in $\widehat{X}$.
In the case of classical solutions, the characteristic curves (2.5) are represented by a linear superposition,
@@
We wish to rewrite $\widehat{X}(x,t)$ in a form with similar linear superpositions, modulo compositions and left-inverses. Note that one can show the family of functions $\{\widehat{X}(\cdot,t)\mid t\in\Omega_{t}\}$ can have a slowly decaying Kolmogorov width, except in special cases with simple initial data $u_{0}$. Consider the stationary shock example shown in Fig. 5. The shock is represented by the constant region $\Upsilon_{t}$ in $\widehat{X}$, which is expanding over time. The jump in the derivative in $\widehat{X}$ travels to the left and the right, which cannot be efficiently represented by the superposition of a few basis functions.
Figure 5: A stationary shock example for the Burgers’ equation. Due to the jump discontinuity in the initial condition $\hat{u}_{0}$ located at $x=0.5$, the shock forms immediately at initial time. This shock is represented as the constant part of the function $\widehat{X}$ in $\Upsilon_{t}$, and the left-inverse of the rarefied characteristic $\widehat{X}^{+}$ has a jump discontinuity there. As time evolves, the characteristics merge into the shock, so $\Upsilon_{t}$ where $\widehat{X}$ is constant correspondingly expands over time. Consequently, the kinks in $\widehat{X}$ at the left and right endpoints of $\Upsilon_{t}$ travels to the left and to the right, respectively. (For a plot of the evolution of the rarefied characteristics, see Fig. 7 below.)
Our key idea is to cut out the portion of the domain where $\widehat{X}_{0}(\cdot,t)$ represents the shock, by adding the characteristic function of $\Upsilon_{t}$ to it. Let us denote the threshold function by $\rho:=\mathbf{1}_{\mathbb{R}_{+}}.$ We call the modified curves the relief characteristic curves, given by
@@
Once the left-inverse is applied to $\widecheck{X}$, the resulting map $\widecheck{X}^{+}:\Omega_{x}\to\widehat{\Omega}_{x}$ is constant at points in $\Upsilon_{t}$, because the term $\rho(t-\lambda(x))$ introduces the constant region and removes the portion of $\widehat{X}$ which lies in $\Upsilon_{t}$ by adding a sufficiently large constant to $\widehat{X}_{0}$ whenever it is evaluted at a point in $\Upsilon_{t}$. See Fig. 6 for an illustration.
@@
In this representation, the time-dependence is in the coefficients of linear combinations within each composition, and the dependence of the coefficients with respect to $t$ is linear, in a similar manner as the classical case. To elaborate, the first linear combination in (2.33) is $t\varphi_{1}-\varphi_{2}$ with two coefficient values $t$ and 1-1. Only the first coefficient depends on $t$ and the dependence is linear.
In the compositional form, the left-inverse $(\cdot)^{+}$ persists, leaving it an implicit form rather than an explicit one. When we consider the discretization of this form of the entropy solution, we will make use of a trick we call the inverse-bias trick that will allow us to avoid any direct evaluation of the left-inverse (Sec. 4.1).
@@
Figure 6: A diagram depicting the relief characteristics $\widecheck{X}$ after shock formation time $t_{S}$ is shown, along with the rarefied characteristics with multivalued inverse $\widehat{X}_{0}$. The relief characteristics use a different representation of the shock compared to the rarefied characteristics $\widehat{X}$, one that uses the addition of a box function $1_{\Upsilon_{t}}$. One observes that left-inverse $\widecheck{X}^{+}$ has a jump at $s(t)$, and this jump is identical to the corresponding one for $\widehat{X}^{+}$.
@@ 3 Low Rank Neural Representation
(3.2) $\begin{bmatrix}\mid&\mid&\mid\\ \phi_{1}(z)&\phi_{2}(z)&\phi_{3}(z)\\ \mid&\mid&\mid\end{bmatrix}\begin{bmatrix}1\\ t\\ 1\end{bmatrix}\quad\in\operatorname{span}\{\phi_{1},\phi_{2},\phi_{3}\}$
where $z=(x,t\varphi_{1}(x)-\varphi_{2}(x))^{T}$ is the output from the previous step. This expression is superposition of a few vectors, if one views the function values $\phi_{i}(z)$ as column vectors of infinite length. In addition, note that the individual the coefficients $1,t,1]^{T}$ has either a linear dependence on $t$ or is constant.
This structure of repeated compositions closely resembles the repeated compositions in a neural network. The component that requires a close scrutiny is the left-inverse $(\cdot)^{+}$ and its treatment will be a key discussion in the coming sections (Sec. 4.1 and Sec. 5.1). For the moment, let us focus on establishing an analogy with feedforward neural networks. We begin by introducing notations and definitions of a neural network. Our notations are similar to those in common use (see, e.g. [22, 3], for similar definitions). Throughout, the superscript $\ell$ will index the layer number.
@@
We will use the shorthand h(W,B):=h(;W,B)h(W,B):=h(\ $\cdot\$;W,B).
@@
[…]
@@
The coefficients $\{\gamma^{\ell}_{i},\theta^{\ell}_{i}\}$ correspond to the coefficients $\{1,t\}$ used in the linear combinations at each compositional step in (3.1). One notes that the latter coefficients depend on the parameter $t$ (the time variable), and the role of this parameter is different from either the input $x$ or the fixed weights and biases $\{U_{i}^{\ell},V_{i}^{\ell}\}$ prescribing the linear spaces the weights and biases belong to. It is therefore natural to view the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ as distinct external parameters; a specific coefficient set $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ correspond to a specific neural network belonging to the family of neural networks whose weights and biases satisfy (3.10). The dependence on $t$ in (3.1) would correspond to the dependence of the coefficients $\{\gamma_{i}^{\ell},\theta_{i}^{\ell}\}$ on $t$. On the other hand, the fixed weights and biases correspond to the functions in the low rank representation that do not depend on $t$, e.g. the functions $\{\varphi_{i},\phi_{j},\psi_{k}\}$ in (3.2).
@@
    (3.14) $\left\{\begin{aligned} U^{\ell}_{i}&=u^{\ell,L}_{i}\bullet u^{\ell,R}_{i},\\ V^{\ell}_{i}&=v^{\ell,L}_{i}\bullet^{\prime}v^{\ell,R}_{i},\end{aligned}\right.\quad\text{ where }\quad\bullet,\bullet^{\prime}\in\{\otimes,{}_{d}\otimes,\otimes_{d}\},$
@@
    coefficient sets $C^{\ell},D^{\ell}\subset\mathbb{R}^{r}$ collected in
@@
In summary, a LRNR is a family of feedforward networks that have weights and bias at each individual layer that belong to a fixed linear subspace of dimension at most $r$. The aim of the following sections is to establish the connection between LRNR (Def. 3.3) and the entropy solution (3.1). Throughout this work, our focus is on scalar problems in a single spatial dimension, so the input and output dimensions are one, that is, $M_{0}=M_{L}=1$.